OpenStudy (anonymous):

Calculate the heat needed at 0 Celsius to make each of the following changes of state. Indicate whether heat was absorbed or released. a. calories to melt 65g of ice b. calories to melt 17g of ice c kilocalories to freeze 225 g of water

OpenStudy (anonymous):

You need to use the specific heat of water. The formula for heat is Heat = (Mass [g]) * (Change in Temperature) * (Specific Heat [capacity]. The problem is a little difficult because the change in temperature isn't there. We'll just use 1. a. (65)*(1)*(4.186) = ~270 calories added b. (17)*(1)*(4.186) = ~70 calories c. [(225)*(1)*(4.186)]/1000 = ~.90 kilocalories removed.

OpenStudy (anonymous):

I found this article earlier today which is sorta pertinent: http://www.wired.com/wiredscience/2011/07/how-much-ice-do-you-need-for-your-drinks/